# Scatterless interferences: Delay of laminar-to-turbulent flow transition by a lattice of subsurface phonons

### v1

@@ Discussion
Each PSub unit is modeled as a finite linear elastic metamaterial consisting of 5 rod unit cells with one local mass-spring resonator in the center of the unit cell [26, 30]. The PSub is free to deform at the edge interfacing with the flow (top) and is fixed at the other end (bottom). Each individual PSub is allowed to deform independently from the adjacent rigid wall and from the motion of neighboring PSubs; its top surface deformation takes a uniform profile across the fluid-PSub interface region [30]. The length of the unit cell along the wall-normal direction is $L_{\mathrm{UC}}=1$ $\mathrm{cm}$ (i.e., total PSub length is $5$ $\mathrm{cm}$). The resonator frequency are set to $\Omega_{\mathrm{res}}=2000$ $\mathrm{Hz}$ by tuning the resonators point masses to be ten times heavier than the total mass of the unit-cell base ($m_{\mathrm{res}}=10\times\rho L_{\mathrm{UC}}$), where $\rho$ is the base material density. Hence, the stiffness of the resonator spring is $k_{\mathrm{res}}=m_{\mathrm{res}}(2\pi f_{\mathrm{res}})^{2}$. The base is composed of ABS polymer with density of $\rho=1200$ $\mathrm{kg/m^{3}}$ and Young’s modulus of $E=3$ $\mathrm{GPa}$. Material damping is modeled as viscous proportional damping with constants $q_{1}=0$ and $q_{2}=6\times 10^{-8}$ [30].
The Navier-Stokes equations are integrated using a time-splitting scheme [34, 35, 36] on a staggered structured grid system. A two-node iso-parametric finite-element model is used for determining the PSub nodal axial displacements, velocities, and accelerations [40] where time integration is implemented simultaneously with the flow simulation using an implicit Newmark algorithm [41]. Since the equations for the fluid and the PSub are inverted separately in the coupled simulations, a conventional serial staggered scheme [42] is implemented to couple the two sets of time integration. More details on the computational models and numerical schemes used are detailed in Refs. [26, 30]. The relative geometric dimensions of the PSubs forming each of the square and hexagonal lattices are to scale as shown in Fig. 4, with more information available in the Supplemntal Materials document.
@@ Acknowledgments
This work utilized the Summit supercomputer, which is supported by the National Science Foundation (awards ACI-1532235 and ACI-1532236), the University of Colorado Boulder, and Colorado State University. The Summit supercomputer is a joint effort of the University of Colorado Boulder and Colorado State University. This work also utilized the Alpine high performance computing resource at the University of Colorado Boulder. Alpine is jointly funded by the University of Colorado Boulder, the University of Colorado Anschutz, and Colorado State University and with support from NSF grants OAC-2201538 and OAC-2322260.

### v2

@@ Discussion
Each PSub unit is modeled as a finite linear elastic metamaterial consisting of 5 rod unit cells with one local mass-spring resonator in the center of the unit cell [26, 30]. The PSub is free to deform at the edge interfacing with the flow (top) and is fixed at the other end (bottom). Each individual PSub is allowed to deform independently from the adjacent rigid wall and from the motion of neighboring PSubs; its top surface deformation takes a uniform profile across the fluid-PSub interface region [30]. The length of the unit cell along the wall-normal direction is $L_{\mathrm{UC}}=1$ $\mathrm{cm}$ (i.e., total PSub length is $5$ $\mathrm{cm}$). The resonator frequency is set to $\Omega_{\mathrm{res}}=2000$ $\mathrm{Hz}$ by tuning the resonator’s point mass to be ten times heavier than the total mass of the unit-cell base ($m_{\mathrm{res}}=10\times\rho L_{\mathrm{UC}}$), where $\rho$ is the base material density. Hence, the stiffness of the resonator spring is $k_{\mathrm{res}}=m_{\mathrm{res}}(2\pi f_{\mathrm{res}})^{2}$. The base is composed of ABS polymer with density of $\rho=1200$ $\mathrm{kg/m^{3}}$ and Young’s modulus of $E=3$ $\mathrm{GPa}$. Material damping is modeled as viscous proportional damping with constants $q_{1}=0$ and $q_{2}=6\times 10^{-8}$ [30].
The Navier-Stokes equations are integrated using a time-splitting scheme [41, 42, 43] on a staggered structured grid system. A two-node iso-parametric finite-element model is used for determining the PSub nodal axial displacements, velocities, and accelerations [47] where time integration is implemented simultaneously with the flow simulation using an implicit Newmark algorithm [48]. Since the equations for the fluid and the PSub are inverted separately in the coupled simulations, a conventional serial staggered scheme [49] is implemented to couple the two sets of time integration. This scheme has been extensively verified, yielding excellent agreement with the experimentally validated linear theory with a maximum deviation of 0.05% in the predicted perturbation energy growth [50].More details on the computational models and numerical schemes used are detailed in Refs. [26, 30]. The relative geometric dimensions of the PSubs forming each of the square and hexagonal lattices are to scale as shown in Fig. 4, with more information available in the Supplemntal Materials document.
@@ Acknowledgments
This research project is currently being supported by Office of Naval Research Multidisciplinary University Research Initiative (MURI) Grant Number N0001421268. The presented work utilized the Summit supercomputer, which is supported by the National Science Foundation (awards ACI-1532235 and ACI-1532236), the University of Colorado Boulder, and Colorado State University. The Summit supercomputer is a joint effort of the University of Colorado Boulder and Colorado State University. This work also utilized the Alpine high performance computing resource at the University of Colorado Boulder. Alpine is jointly funded by the University of Colorado Boulder, the University of Colorado Anschutz, and Colorado State University and with support from NSF grants OAC-2201538 and OAC-2322260.